OpenStudy (anonymous):

PLEASE PLEASE HELP been stuck on this for more then 2 hours! A zero compound bond is a bond that is sold now at a discount and will pay its face value at the time when it matures, no interest payment are made. A zero-coupon bond can be redeemed in 20 years for $10000. How much should you be willing to pay for it now if want a return of. (a) 13% compounded daily. For a return of 13% compounded daily, should pay $_____ (b) 13% compounded continuously? for a return of 13% continuously should pay $_____

OpenStudy (e.mccormick):

You are dealing with two compound interst formulas. First, is \(P = C (1 + \frac{r}{n})^{nt}\) and the second is \(P = C e^{rt}\) where P = future value C = initial deposit r = interest rate (expressed as a fraction: eg. 0.06) n = # of times per year interest is compounded t = number of years invested e = The constant e. In your case, they give you P=10,000, t=20, r=13%=0.13, and n=365 (well, that is roughlty dayly, there is also leap years so... not sure if they care about that.)

OpenStudy (e.mccormick):

So you plug in what you know and solve for C, what you would pay for it now.

OpenStudy (anonymous):

I got 13457.51 is that right?

OpenStudy (e.mccormick):

You should get something LESS than the end value. Basically yhey are saying, if you pay C, then C+compounded interest=10,000.

OpenStudy (e.mccormick):

a) is: \[P = C (1 + \frac{r}{n})^{nt} \Rightarrow 10,000 = C (1 + \frac{.13}{365})^{(365)(20)}\] b) is: \[P = C e^{rt} \Rightarrow 10,000 = C e^{(.13)(20)}\]

OpenStudy (anonymous):

\[10000(1+\frac{ 0.13 }{ 365 })^{7300}\] this is what I plug in

OpenStudy (anonymous):

Oh trying to find C

OpenStudy (e.mccormick):

Yah, I guessed where your mistake was. The concept is that they are giving you the output and asking you to find the input.

OpenStudy (anonymous):

Thats what I been doing wrong the whole Time!! do you do the log to get rid of 7300? well how do you get rid of 7300

OpenStudy (e.mccormick):

No need to get rid of it. Just use it. \[10000=C (1 + \frac{.13}{365})^{(365)(20)}\Rightarrow \frac{10000}{(1 + \frac{.13}{365})^{7300}}=C\]

OpenStudy (e.mccormick):

So about $743.08. Not much to get $10,000. But with 20 years of interest, that does not surprise me.

OpenStudy (anonymous):

I just get 10000

OpenStudy (e.mccormick):

You need to solve for C, so you divide both sides by the known part that is normally multiplying C. That is all I did.

OpenStudy (anonymous):

yeah i see what I did wrong!

OpenStudy (e.mccormick):

Now, do the same sort of thing to the pert one. What do you get there?

OpenStudy (anonymous):

I got 742.74

OpenStudy (e.mccormick):

Same as me! I think all that happened was that after using principal to find interest or how much you get after so long, they tossed one at you with, "Here is the answer, find the question!" and it caught you off guard.

OpenStudy (anonymous):

lol thank you soo much! was stuck on this question for more then 2 hours an no one helped! again thank you!!!

OpenStudy (e.mccormick):

Yah, I saw it earlier. I just took a break in my Linerar homework and saw it was still there!

OpenStudy (anonymous):

still thank you for coming back for it! ^_^

OpenStudy (e.mccormick):

np. Don't have too much fun!

OpenStudy (anonymous):

haha try not to!